OpenStudy (anonymous):

Can yall please help find the inverse of y=5x^2 -8? Thanks!!

OpenStudy (anonymous):

Do you know the first step in finding the inverse of a function?

OpenStudy (anonymous):

Hint: It has to do with the x and the y

OpenStudy (alexandervonhumboldt2):

x=5y^2-8 is the rirst step

OpenStudy (anonymous):

switching the x and y

OpenStudy (alexandervonhumboldt2):

(x+8)/5=y^2

OpenStudy (alexandervonhumboldt2):

y=sqrt((x+8)/5)

OpenStudy (alexandervonhumboldt2):

@gabby13

OpenStudy (anonymous):

im getting confused after adding the 8 to both sides... what do I do next?

OpenStudy (alexandervonhumboldt2):

@Loser66

OpenStudy (alexandervonhumboldt2):

@SolomonZelman

OpenStudy (loser66):

@AlexandervonHumboldt2 when someone is trying to help you, please, respect, don't tag others.

OpenStudy (anonymous):

can someone help me?

OpenStudy (anonymous):

Lets see if I can help \[y=5x^2-8\] \[x=5y^2-8\] \[x+8=5y^2-8+8\] \[(x+8)/5=(5y^2)/5\] \[(x+8)/5=y^2\] and to cancel the power of 2 on the y, we use the square root function \[sqrt((x+8)/5)=sqrt(y^2)\] \[sqrt((x+8)/5)=y\]

OpenStudy (anonymous):

Thank you!!! You helped me so much!!!